solve the system of linear equations using the guass-jordan method.
x-2y=8
3x+4y=4

Answer by ks(2)

x-2y=8 ------------ (1)
3x+4y=4 ----------- (2)
Multiply equation (1) by 2
2x-4y=16 ---------(3)
3x+4y=4 ---------- (4)
Subtract equation (4) from (3)
-x=12
x=-12
Substitute x=-12 in(1)
-12-2y=8
-12-8=2y
2y=-20
y=-10
Hence , x=-12, y=-10